jhonyy9 (jhonyy9):

- something incredible - how is this possible ? 1=2 ???????????

jhonyy9 (jhonyy9):

anonymous Best Response Medals 0 ln(x)=∫1/x=∫2/(2x)=ln(2x) 3 months agoReport Abuse 46 anonymous Best Response Medals 0 hi @ganeshie8 3 months agoReport Abuse qh 99 ganeshie8 Best Response Medals 0 im familiar with this il let somebody else figure this out this time :) 3 months agoReport Abuse 46 anonymous Best Response Medals 0 how do you think can i have new conjecture in world of primes ? because i think that i have got some - 1-2-3 ??? 3 months agoReport Abuse qh 99 Kainui Best Response Medals 1 Ahhh yeah this is a fun trick to fool people I forgot about this one haha. 3 months agoReport Abuse 46 anonymous Best Response Medals 0 @ganeshie8 excuze me for your derange - i m technician electronics - the mathematics is one my bige hobby - and i think that i havent with who i can talking these interestingly math idea what i got ,i have - you dont have something idea please for how can i share these with math-world - yes if it's worth 3 months agoReport Abuse qh 99 Kainui Best Response Medals 1 Here's two hints: Rules of logarithms Indefinite integrals 3 months agoReport Abuse 46 anonymous Best Response Medals 0 @ganeshie8 and because you are here @Kainui how you think it please ? can we name the twins with sum of members +/- 1 generate in first round new twins too - so can we name these twins perfectly twins ? opinions ? - this may be my first conjecture - that is possibly 3 months agoReport Abuse 46 anonymous Best Response Medals 0 for example 5-7 5+7=12 12 +/- 1 = 11-13 so this mean that 5-7 are perfectly twins 3 months agoReport Abuse 46 anonymous Best Response Medals 0 29-31 29+31 = 60 60+/-1 = 59-61 so these same first example 3 months agoReport Abuse 46 anonymous Best Response Medals 0 29-31 are perfectly twins 3 months agoReport Abuse qh 99 Kainui Best Response Medals 1 Oh that's really cool I like that! 3 months agoReport Abuse 46 anonymous Best Response Medals 0 and what is your opinions about absolutely twins ? so this mean that a perfectly twin like 5-7 with sum of members = 12+/-1 generate new twin like in this case 11-13 and than we assum 11+13=24 but 24x3=72+/-1 = 71-73 so this mean that 5-7 are perfectly twins 3 months agoReport Abuse 46 anonymous Best Response Medals 0 hope so much that i write it understandably - pardon for my grammatically errors 3 months agoReport Abuse 46 anonymous Best Response Medals 0 @ganeshie8 so this is the way how i have got these biger twins of 2591-2593 3 months agoReport Abuse qh 99 Kainui Best Response Medals 1 Hmm I sort of understand, but I don't know what absolutely twins are. Does this mean it leads to a chain of twin primes? Specifically what is this? 24x3=72 why 3? 3 months agoReport Abuse 46 anonymous Best Response Medals 0 yes 72+/-1 = 71-73 new twin 3 months agoReport Abuse 46 anonymous Best Response Medals 0 this is a way what i have got to generate new primes and /or twins 3 months agoReport Abuse qh 99 Kainui Best Response Medals 1 What is 3, why not 4 or 5? 3 months agoReport Abuse qh 99 Kainui Best Response Medals 1 24x5=120 but 119 and 121 are not twin primes. 3 months agoReport Abuse 46 anonymous Best Response Medals 0 and again something interestingly 2+2=4+/-1 = 5-7 5+7=12+/-1=11-13 11+13=24x3=72+/-1=71-73 71+73=144 144x3=432+/-1=431-433 431+433=864 864x3=2592+/-1 = 2591-2593 3 months agoReport Abuse 46 anonymous Best Response Medals 0 this is the biger directly family tree of twins what i have got using this way 3 months agoReport Abuse 46 anonymous Best Response Medals 0 @ganeshie8 - how you like it - please ? 3 months agoReport Abuse qh 99 Kainui Best Response Medals 1 Hmm it seems kind of arbitrary and doesn't hit all the twin primes but it is interesting and fun to play with. 3 months agoReport Abuse 46 anonymous Best Response Medals 0 yes this is true but remember in case of human family tree not everybody are kin or direct descendant - yes ? 3 months agoReport Abuse 46 anonymous Best Response Medals 0 so why not can being this possibly so in the world of primes too ? 3 months agoReport Abuse 46 anonymous Best Response Medals 0 @ganeshie8 @ParthKohli ????- opinion 3 months agoReport Abuse qh 99 Kainui Best Response Medals 1 True true, not all primes are related. This is interesting I'm gonna keep playing with these and I'll tell you if I find anything. Right now just working on making a program to find the longest chain of these in Java. Do you program? 3 months agoReport Abuse 46 anonymous Best Response Medals 0 sorry no - than you read my above wrote words i m technician electronist and the mathematics is one of my big big hobby - i like -love it because my opinion that is the best logically subject from all over the world 3 months agoReport Abuse 46 anonymous Best Response Medals 0 these i have wrote to @ganeshie8 to and than i have something interestingly like new conjecture - for example - so i dont know how or with who need talking this for i can deciding than is right or not is importantly or not ,---- ??? 3 months agoReport Abuse qh 99 Kainui Best Response Medals 1 Well you should spend some time learning just a little bit of basic programming, it would be a great tool for you! 3 months agoReport Abuse 46 anonymous Best Response Medals 0 why do you think it so ? i know that in the computer world use to much prims and ... 3 months agoReport Abuse 46 anonymous Best Response Medals 0 and so in the bank sector with money to generate different codes of bankcards or bank accounts - i have heard it so - and because have said me one my college of work and i have got these two prims from what was created the cvv code of my bank card - so this mean that this make sens - right 3 months agoReport Abuse 61 matlee Best Response Medals 0 I dont get it but if it could be explained i wanna see 3 months agoReport Abuse 99 SolomonZelman Best Response Medals 1 ln(x)=∫dxx=∫22xdx=ln(2x)+C=ln(x)+ln(2)+C=ln(x)+C 3 months agoReport Abuse 99 SolomonZelman Best Response Medals 1 This is why, also, ddxln(αx)=ln(x)+ln(α)=1x+0=1x (and using the chain rule, the result obained is same) 3 months agoReport Abuse qh 99 Kainui Best Response Medals 1 One thing to keep in mind is that we can write ALL numbers of these form (mod 6 arithmetic): 6n, 6n+1, 6n+2, 6n+3, 6n+4, 6n+5 All of these for n>0 are composite and can be factored except 6n+1 and 6n+5 which is really the same as 6n+1 and 6n-1. All twin primes greater than the pair (3,5) are of the form 6n-1 and 6n+1 so if you add them together like you've done: 6n-1 + 6n -1 = 12n = 6(2n) then you have a possibility of having 6(2n)-1 and 6(2n)+1 for another twin prime again! 3 months agoReport Abuse 46 anonymous Best Response Medals 0 @Kainui - thank you for your words but little missed term there so 6n-1+6n-1= 12n-2 and not just 12n like how you wrote above 3 months agoReport Abuse 46 anonymous Best Response Medals 0 i think you wan adding there 6n-1+6n+1 = 12n - in this way is right sure 3 months agoReport Abuse 46 anonymous Best Response Medals 0 @saifoo.khan @satellite73 please your opinion about these ? thank you very much 3 months agoReport Abuse 46 anonymous Best Response Medals 0 @jim_thompson5910 - please 3 months agoReport Abuse 46 anonymous Best Response Medals 0 @SamBot - please read these comments and i wait your opinion about them 3 months agoReport Abuse 40 SamBot Best Response Medals 0 um im only in middle school i have no idea!! 3 months agoReport Abuse 46 anonymous Best Response Medals 0 Dear CHAMPIONS ! - samthing new bout primes 5 - 7 ------ 5+7=12 +/- 1 = 11 - 13 -------- 11+13=24 x 3 = 72 +/- 1 = 71 - 73 -------- 72+73=144 x 3 = 432 +/- 1 = 431 --- 433 ------------ 431+433=864 x 3 = 2592 +/- 1 = 2591 ----- 2593 ---------------- - so than you see these all twins what have resulted begin this line with 5 - 7 every resulted twins has members ending in 1 and 3 - so not is this very very interesting ? - than we begin it by 17 --- 19 ---------- 17+19=36 x3 = 108+/- 1 = 107 --- 109 ------------ 107 +109 = 216 x3 = 648+/-1 = 647 --- 649 ------------ - so than you see them started by 17 and 19 every resulted twins has mambers ended in 7 and 9 - not is interesting too ? 29 --- 31 --------- 29+31 = 60 +/-1 = 59 --- 61 --------- 59+61=120 x3 = 360+/-1= 359 --- 361 ------------ - so every two twins has members ending in 59 and 61 - how much interesting it these ? 3 months agoReport Abuse 77 FaiqRaees Best Response Medals 0 Originally Posted by @SolomonZelman ln(x)=∫dxx=∫22xdx=ln(2x)+C=ln(x)+ln(2)+C=ln(x)+C End of Quote In this I think you've done a mistake. ln(2)≠0 2 months agoReport Abuse 71 baru Best Response Medals 0 He's lumping constants together. ln(2) + C is another constant 2 months agoReport Abuse 46 anonymous Best Response Medals 0 @jim_thompson5910 please than you check the above wrote case about super twins so than what you think about star twins please ? thank you in advance 19 days agoReport Abuse Type your reply